# PV-S3111Photovoltaic-Semi Supervised Segmentation : Advancing Automatic Photovoltaic Defect Detection using Semi-Supervised Semantic Segmentation of Electroluminescence Images

### v3

@@ 4.2 Discussion and Analysis
Figure 10: Comparison of confusion matrix for supervised and semi-supervised learning: The image on the left shows the confusion matrix for supervised learning with 20% labeled images and the right image shows semi-supervised learning with 20% labeled images. The x-axis represents the predicted class labels and y axis represents the true class labels. The values are shown in the log scale for better visualization.
As the name suggests, using a confusion matrix we can identify the classes that are confusing for the model. We compare supervised and semi-supervised learning models’ performance using a confusion matrix (Figure 10). The supervised model misclassifies many pixels, resulting in low performance. In contrast, the semi-supervised model shows fewer misclassifications, leading to improved defect detection. The class imbalance as evident from Figure  4 introduced some bias in the predictions. Specifically, the semi-supervised model correctly classifies significantly more pixels for each defect class compared to the supervised model, indicating the effectiveness of the PV-S3 model. We observe that semi-supervised learning correctly classified 387157 (Contact), 48175 (Interconnect), 2582157 (Crack), and 53188 (Corrosion) pixels compared to the baseline (supervised) model which classified only 48050 (Contact), 1070 (Interconnect), 757701 (Crack), and 33429 (Corrosion) pixels correctly, which shows a significant improvement for PV-S3 model. However, the semi-supervised model struggles most with the contact defect class. This is due to its reliance on limited labeled data, hindering effective error propagation and feature learning. The confusion matrix in Figure 10 visually represents these findings with logarithmic values for clarity.

### v4

@@ 4.2 Discussion and Analysis
Figure 10: Comparison of confusion matrix for supervised and semi-supervised learning: The image on the left shows the confusion matrix for supervised learning with 20% labeled images and the right image shows semi-supervised learning with 20% labeled images. The x-axis represents the predicted class labels and y axis represents the true class labels. The pixel values in the confusion matrices are transformed using natural logarithmic scale (base e) to enhance visibility and interpretation, especially due to the significant class imbalance in the dataset. The defects are abbreviated as CT: Contact, IC: Interconnect, CR: Crack and CO: Corrosion
We compare the supervised and semi-supervised learning models using a confusion matrix (Figure 10), which uses a $log_{e}$ scale to enhance visual clarity. The supervised model misclassifies many pixels, resulting in low performance. In contrast, the semi-supervised model shows fewer misclassifications, leading to improved defect detection. The class imbalance, as shown in Figure 4, introduces some bias in the predictions. Specifically, the semi-supervised model correctly classifies a higher number of pixels for each defect class compared to the supervised model, demonstrating the effectiveness of the PV-S3 model. To provide quantitative insight, we report both the absolute number of correctly classified pixels and their natural logarithm (loge) values, as shown in the confusion matrix. For the semi-supervised model, the correctly classified pixel counts are: $12.87$ ($\log_{e}(387,157)$) for Contact, $10.78$ ($\log_{e}(48,175)$) for Interconnect, $14.76$ ($\log_{e}(2,582,157)$)) for Crack, and $10.88$ ($\log_{e}(53,188)$)) for Corrosion. In comparison, the supervised model correctly classifies $10.78$ ($\log_{e}(48,050)$) for Contact, $6.97$ ($\log_{e}(1,070)$) for Interconnect, $13.54$($\log_{e}(757,701)$) for Crack, and $10.42$ ($\log_{e}(33,429)$) for Corrosion. These numbers illustrate a clear improvement by the PV-S3 model, especially for low-frequency classes.
Despite being the most frequent defect in terms of both image count and pixel area (Figure 4), the semi-supervised model performs poorly on the Contact class. This counterintuitive result stems from the physical interplay of degradation processes in PV modules. Contact failures, such as corrosion or front grid interruptions, often co-occur with other defects like Cracks and Interconnect failures due to shared stress and failure pathways. For instance, corrosion near solder pads can induce local stress that leads to cracking, while broken interconnects can cause localized heating that accelerates corrosion. These defects frequently overlap spatially and exhibit similar visual textures, complicating the model’s ability to distinguish among them—especially under limited supervision. The confusion matrix in Figure 10 highlights this challenge, showing that Contact pixels are often misclassified as Cracks. Even prior works using fully supervised models [18] trained on the same dataset achieve only 66% precision and 51% recall on the Contact class. This consistent difficulty across settings suggests that the issue is not merely due to label scarcity but is intrinsic to the complex, interacting nature of defects in real-world PV cells.